# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Choices: [(a) Sarah couldn’t believe her eyes. On the table were plates full of blueberry pancakes, blueberry muffins, and even blueberry sauces., (b) Order our blueberry pancakes today! They’re always hot and delicious., (c) As the blueberry fruit ripens, it turns from green to purple to dark blue.]
Answer index: 2
Subject: Language Arts
Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-2
Description: Review the passage.
I thought the matter over during the next day, Sunday, and finally resolved upon the third day of September, as the day upon which I would make a second attempt to secure my freedom. I now had three weeks during which to prepare for my journey.
From Frederick Douglass, Narrative of the Life of Frederick Douglass
Look at the text in **bold**. What hidden reality is shown through this text?
Choices: [(a) It shows that although Douglass’s friends behaved kindly, they secretly envied him., (b) It shows that although the shipyard was calm, Douglass experienced tension at home., (c) It shows that although the city felt peaceful, enslaved people were preparing to rebel., (d) It shows that although Douglass’s plan was working, he still had worries and doubts.]
Answer index: 3
Subject: Language Arts
Skill: are-the-modifiers-used-correctly
Description: Which sentence is correct?
@@
Choices: [(a) **Stuck in traffic**, Hakim grew drowsy and time slowed to a standstill; eventually, he decided to pull over and take a nap., (b) **Stuck in traffic**, time slowed to a standstill and Hakim grew drowsy; eventually, he decided to pull over and take a nap.]
Answer index: 0
Subject: Language Arts
Skill: recall-the-source-of-an-allusion
Description: What is the source of the allusion in the sentence below?
The headline read: "Smart Phones Becoming **Big Brother**."
Choices: [(a) literature, (b) a song]
Answer index: 0
Subject: Language Arts
Skill: transitions-with-conjunctive-adverbs
Description: Complete the text with the better conjunctive adverb.
@@
In China, mountainous areas are being flattened to make more room for new construction; ___, in the United States, mountaintops are being removed so that miners can access coal reserves.
Choices: [(a) therefore, (b) likewise]
Answer index: 1
Subject: Language Arts
Skill: determine-the-authors-point-of-view
Description: Read the text.
When I decided it was time for a new hobby, I signed up for a "Clowning for Fun and Profit" class at my local arts center. I cheerily skipped through the door on the first day of class, but my heart sank as soon as I spotted the instructor, Bobo. Bobo turned out to be a sad-faced clown who refused to speak. It was clear he planned to teach entirely in pantomime. As he silently motioned for me to find my seat, I resigned myself to a bad experience.
@@
So imagine my surprise when this crazy class turned out to be invaluable to me! Although I didn’t hear a single spoken word for five hours, Bobo somehow managed to express his ideas clearly and in an engaging way. By the end of the first session, he had expertly taught me how to make balloon animals, apply makeup, juggle beanbags, and develop my own personal clown character. As I eagerly await the second session, "All About Unicycles," I will begin clowning around as a volunteer at a local children’s hospital.
Which sentence best summarizes the author’s point of view about the clown class?
Choices: [(a) Bobo uses eccentric teaching methods to efficiently teach this course, making him an excellent clowning instructor., (b) Bobo’s sad demeanor and poor pantomime skills make it difficult for his students to learn effective clowning techniques.]
Answer index: 0
Subject: Language Arts
Skill: compare-two-texts-with-different-genres
Description: Read the following text from a diary. This excerpt is about life for Jews in Europe at the beginning of World War II.
@@
After May 1940 good times rapidly fled: first the war, then the capitulation, followed by the arrival of the Germans, which is when the sufferings of us Jews really began. Anti-Jewish decrees followed each other in quick succession. Jews must wear a yellow star, Jews must hand in their bicycles, Jews are banned from trams and are forbidden to drive. Jews are only allowed to do their shopping between three and five o’clock and then only in shops which bear the placard "Jewish shop." Jews must be indoors by eight o’clock and cannot even sit in their own gardens after that hour. Jews are forbidden to visit theaters, cinemas, and other places of entertainment.
From Anne Frank, "Anne Frank: The Diary of a Young Girl". Copyright 1952 by Otto H. Frank
Based on this excerpt, what can you conclude about life for Jews in Europe at the beginning of World War II?
Choices: [(a) Jewish people could no longer afford to buy theater tickets., (b) Jewish people could not associate with one another., (c) Jewish people had many of their rights taken away.]
Answer index: 2
@@
Subject: Language Arts
Skill: read-and-understand-informational-passages
Description: Review the **seventh paragraph**.
title: Meditating in School
@@
This is where the need for good, sound research comes in. Past studies showing that mindfulness helps medical patients are not enough. To really determine whether mindfulness helps improve grades, research needs to focus specifically on mindfulness in schools—and what impact, if any, it has. That way, educators can make informed decisions about where to put their resources. Until then, whether mindfulness is included in classroom programming will remain the choice of individual teachers.
Based on the text, what still needs to be determined about mindfulness in schools?
Choices: [(a) whether meditation affects different students in different ways, (b) how to raise the funds needed to teach meditation in schools, (c) whether mindfulness meditation has a positive impact in a school setting]
Answer index: 2
Subject: Language Arts
Skill: choose-evidence-to-support-a-claim
Description: Read the claim below.
Gardeners should never use pesticides on their plants.
Select the piece of evidence that best supports this claim.
@@
Choices: [(a) Organic pesticides tend to cost more than conventional ones, and they may be less effective., (b) Pets and children often play in and around gardens, where they can be exposed to pesticides.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-the-giver-part-1
Description: Reread the following passage from The Giver. In this excerpt, Jonas, the main character, recalls a time when he saw an unidentified aircraft.
@@
Frightened was the way [Jonas] had felt a year ago when an unidentified aircraft had overflown the community twice. He had seen it both times. Squinting toward the sky, he had seen the sleek jet, almost a blur at its high speed, go past, and a second later heard the blast of sound that followed. Then one more time, a moment later, from the opposite direction, the same plane.
From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry
Look at the text in **bold**. How does this paragraph contribute to the meaning or tone of the passage?
Choices: [(a) It creates an unsettling tone., (b) It creates a ridiculous tone., (c) It creates a playful tone., (d) It creates a serene tone.]
Answer index: 0
Subject: Language Arts
Skill: use-personification
Description: Complete the sentence so that it uses personification.
The sorceress peered into the mirror, noticing how the amulet ___ her neck.
Choices: [(a) barely clasped around, (b) gently embraced]
Answer index: 1
Subject: Language Arts
Skill: identify-plagiarism
Description: Compare the student text with the source.
Source: Carolyn Cinami DeCristofano, "A Black Hole Is Not a Hole", page 5. Published by Charlesbridge, 2012:
Nothing can out-tug a black hole. No army of tow trucks, no convoy of supersized earth haulers, no fleet of giant rocket engines.
Student text:
The pull of a black hole is so strong that nothing can pull stronger, not even a "fleet of giant rocket engines" (DeCristofano 5).
Is the student text plagiarized?
Choices: [(a) No, it is not plagiarized., (b) Yes, because it fails to use quotation marks.]
Answer index: 0
Subject: Language Arts
Skill: identify-counterclaims
Description: Consider this claim:
Poems that rhyme are better than poems that do not rhyme.
Now consider how someone might argue against this. Which sentence presents a counterclaim to the above claim?
Choices: [(a) "Beowulf" is one of the oldest and longest poems in existence., (b) Poetry that doesn’t rhyme allows for more freedom of expression.]
Answer index: 1
Subject: Language Arts
Skill: trace-an-argument
Description: Read the text.
title: Early to Bed and Early to Rise
@@
Whether you’re naturally a night owl or an early bird is mostly based on your body’s chronotype, which Merriam-Webster’s online dictionary defines as "the internal circadian rhythm or body clock of an individual that influences the cycle of sleep and activity in a 24-hour period." In short, we are born wired one way or the other. However, it is possible to change your habits. And given the benefits of being an early riser, it may be worth the effort.
What is the author’s main claim or argument?
Choices: [(a) There are just as many benefits to being a night owl as an early riser., (b) It’s important to change how society perceives night owls., (c) It is better to be an early riser than a night owl.]
Answer index: 2
Subject: Language Arts
Skill: identify-active-and-passive-voice
Description: Which sentence is in passive voice?
Choices: [(a) On April 18, 1906, an earthquake and its fiery aftermath devastated the city of San Francisco., (b) On April 18, 1906, San Francisco was devastated by an earthquake and its fiery aftermath.]
Answer index: 1
Subject: Language Arts
Skill: is-the-pronoun-reflexive-or-intensive
Description: Is the pronoun in bold reflexive or intensive?
@@
Bryce uses an elaborate system of notes to remind **himself** of what he has left to do.
Choices: [(a) intensive, (b) reflexive]
Answer index: 1
Subject: Language Arts
Skill: interpret-figures-of-speech
Description: What does the **metaphor** in this text suggest?
When Bobby lost his job, he was devastated. The only light in the sea of darkness was the prospect of pursuing a new career.
Choices: [(a) There was a benefit to Bobby’s job loss., (b) Having to pursue a new career was the worst part of Bobby’s job loss.]
Answer index: 0
Subject: Language Arts
Skill: identify-the-verb-mood
Description: Which verb mood is used in the **bold** part of the sentence?
To help maintain your energy and focus throughout the day, **eat a healthy breakfast**.
Choices: [(a) imperative, (b) subjunctive]
Answer index: 0
Subject: Language Arts
Skill: classify-figures-of-speech
Description: Which figure of speech is used in this text?
The water made a sound **like kittens lapping**.
—Marjorie Kinnan Rawlings, "The Yearling"
Choices: [(a) metaphor, (b) simile]
Answer index: 1
Subject: Language Arts
Skill: vocabulary-review-trace-an-argument
Description: Look at the word **entrenched** as used in the passage "The Most Important Meal of the Day?"
@@
An Australian study published in 2019 collected health data for over twenty-eight years from multiple countries. The results showed few differences between breakfast skippers and breakfast eaters. The study concluded that "breakfast is not the most important time of the day to eat, even though that belief is really **entrenched** in our society and around the world."
What does **entrenched** mean?
Choices: [(a) highly debatable, (b) quite rare, (c) firmly established]
Answer index: 2
Subject: Language Arts
Skill: analyze-passages-from-the-outsiders-part-2
Description: Review the passage.
@@
Tim Shepard and company were already waiting when we arrived at the vacant lot, along with a gang from Brumly, one of the suburbs. Tim was a lean, catlike eighteen-year-old who looked like the model JD you see in movies and magazines. He had the right curly black hair, smoldering dark eyes, and a long scar from temple to chin where a tramp had belted him with a broken pop bottle. He had a tough, hard look to him, and his nose had been broken twice. Like Dally’s, his smile was grim and bitter. He was one of those who enjoy being a hood. The rest of his bunch were the same way. The boys from Brumly, too. Young hoods—who would grow up to be old hoods. I’d never thought about it before, but they’d just get worse as they got older, not better. I looked at Darry. He wasn’t going to be any hood when he got old. He was going to get somewhere. Living the way we do would only make him more determined to get somewhere. That’s why he’s better than the rest of us, I thought. He’s going somewhere. And I was going to be like him. I wasn’t going to live in a lousy neighborhood all my life.
From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton
Look at the text in **bold**. What does this figurative language suggest?
Choices: [(a) Tim fights for survival every day., (b) Tim has always been cruel to Ponyboy., (c) Tim often does not have enough to eat., (d) Tim attempts to stay hidden at all costs.]
Answer index: 0
@@
Subject: Language Arts
Skill: identify-the-authors-purpose
Description: Review the text.
Good study habits train your brain for success. Applying effective learning methods will not only improve your grades but, more importantly, increase your long-term knowledge.
Cramming at the last minute is ineffective. Instead, try to create a study routine. Schedule specific blocks of time to study every day. Next, find the best environment that helps you concentrate. Some students do better with the silence of a library, while others need the background noise or music of a public space.
Now that you have a time and a place, you need a strategy. Look at your tasks and make a list. Setting small, manageable goals for your study session will help you work efficiently. Do the most difficult assignments first, while your mind is fresh. Finally, do not be afraid to take limited breaks to restore your motivation.
Which best shows that the author’s purpose is "to explain some productive study techniques to readers"?
Choices: [(a) The author argues that students need to work in a place with "background noise" in order to study efficiently., (b) The author outlines a series of best practices to follow to improve study habits.]
Answer index: 1
Subject: Language Arts
Skill: is-it-a-direct-object-or-an-indirect-object
Description: Is the phrase in bold a direct object or an indirect object?
A guard denied **the tourists** entrance to the palace.
Choices: [(a) direct object, (b) indirect object]
Answer index: 1
Subject: Language Arts
Skill: decide-whether-ellipses-are-used-appropriately
Description: Read this text.
On May 25, 1977, "Star Wars" opened in theaters and wowed audiences with groundbreaking special effects.
Is this an appropriate use of an ellipsis?
Choices: [(a) yes, (b) no, because the text with an ellipsis is no longer grammatically correct]
Answer index: 1
Subject: Language Arts
Skill: compound-subjects-and-objects-with-pronouns
Description: Select the correct pronoun to complete the sentence.
Mr. Joyce rarely worries about ___ and Farid, because they’re both so responsible.
Choices: [(a) she, (b) her]
Answer index: 1
Subject: Language Arts
Skill: identify-vague-pronoun-references
Description: Which of the following contains a vague pronoun reference?
Choices: [(a) Bella and her sister posed for a photograph, but when the flash went off, she blinked., (b) Bella and her sister posed for a photograph, but when the flash went off, Bella blinked.]
Answer index: 0
Subject: Language Arts
Skill: what-does-the-punctuation-suggest
Description: What does the following sentence suggest?
@@
Ferries depart from Battery Park in New York and Liberty State Park in New Jersey for the museum on Ellis Island, which commemorates the millions of immigrants who entered the United States through Ellis Island.
Choices: [(a) There is only one museum on Ellis Island., (b) There is more than one museum on Ellis Island.]
Answer index: 0
Subject: Language Arts
Skill: suggest-appropriate-revisions
Description: Read the following text from a student essay. How can the writer best improve his or her **word choice**?
@@
I saw a movie called "Blob Invasion". It was really good. It was about a weird guy who was into science. He worked on bugs and small animals. He made a machine that did things to the bugs and the animals. They became big weird things. One day he transformed himself by accident. He became a big blob! He went around doing bad things to the city. In the end, a boy figured out how to change the blob back into a man. I would recommend this movie to everyone who likes fun science movies.
Choices: [(a) by reducing repetitive language, (b) by using more specific language]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-the-giver-part-2
Description: Review the passage.
"Father? Mother?" Jonas asked tentatively after the evening meal. "I have a question I want to ask you."
From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry
Based on the passage, how does Jonas most likely feel about his family?
@@
Choices: [(a) He finds comfort in the regimented order provided by his family., (b) He wonders why his family’s beliefs are different from the community’s., (c) He feels grateful to have a family that openly discusses tough issues., (d) He is beginning to question his parents and the community.]
Answer index: 3
Subject: Language Arts
Skill: identify-thesis-statements
Description: Which is a thesis statement?
@@
Choices: [(a) School-wide Wi-Fi access is critical to building academic success because it increases communication, allows for more interactive lessons, and expands research opportunities., (b) What effect does having Wi-Fi access in public school classrooms and libraries have on how students perform academically and how well they pay attention in class?]
Answer index: 0
Subject: Language Arts
Skill: classify-logical-fallacies
Description: Which logical fallacy is used in the text?
@@
In this election, we’ve seen the highest voter turnout ever recorded. If you have any doubts about Ryan Mueller’s qualifications, just look at how many people have come out to vote for him.
Choices: [(a) bandwagon fallacy: the assumption that the popular choice is automatically correct, (b) ad hominem: a personal attack against one’s opponent]
Answer index: 0
Subject: Language Arts
Skill: analyze-passages-from-the-outsiders-part-1
Description: Read the following passage from "The Outsiders". In this excerpt, Ponyboy, the narrator, and his friend, another greaser named Two-Bit, are walking with two girls from a rival group called the Socs.

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: choose-the-text-that-matches-the-writers-purpose Description: You want to **inform** someone **about blueberries**. "Inform" means "teach someone about something". What should you write? Choices: [(a) Sarah couldn’t believe her eyes. On the table were plates full of blueberry pancakes, blueberry muffins, and even blueberry sauces., (b) Order our blueberry pancakes today! They’re always hot and delicious., (c) As the blueberry fruit ripens, it turns from green to purple to dark blue.] Answer index: 2
Subject: Language Arts Skill: analyze-passages-from-narrative-of-the-life-of-frederick-douglass-part-2 Description: Review the passage. I thought the matter over during the next day, Sunday, and finally resolved upon the third day of September, as the day upon which I would make a second attempt to secure my freedom. I now had three weeks during which to prepare for my journey. From Frederick Douglass, Narrative of the Life of Frederick Douglass Look at the text in **bold**. What hidden reality is shown through this text? Choices: [(a) It shows that although Douglass’s friends behaved kindly, they secretly envied him., (b) It shows that although the shipyard was calm, Douglass experienced tension at home., (c) It shows that although the city felt peaceful, enslaved people were preparing to rebel., (d) It shows that although Douglass’s plan was working, he still had worries and doubts.] Answer index: 3
@@
Subject: Language Arts Skill: are-the-modifiers-used-correctly Description: Which sentence is correct? Choices: [(a) **Stuck in traffic**, Hakim grew drowsy and time slowed to a standstill; eventually, he decided to pull over and take a nap., (b) **Stuck in traffic**, time slowed to a standstill and Hakim grew drowsy; eventually, he decided to pull over and take a nap.] Answer index: 0
Subject: Language Arts Skill: recall-the-source-of-an-allusion Description: What is the source of the allusion in the sentence below? The headline read: "Smart Phones Becoming **Big Brother**." Choices: [(a) literature, (b) a song] Answer index: 0
@@
Subject: Language Arts Skill: transitions-with-conjunctive-adverbs Description: Complete the text with the better conjunctive adverb. In China, mountainous areas are being flattened to make more room for new construction; ___, in the United States, mountaintops are being removed so that miners can access coal reserves. Choices: [(a) therefore, (b) likewise] Answer index: 1
@@
Subject: Language Arts Skill: determine-the-authors-point-of-view Description: Read the text. When I decided it was time for a new hobby, I signed up for a "Clowning for Fun and Profit" class at my local arts center. I cheerily skipped through the door on the first day of class, but my heart sank as soon as I spotted the instructor, Bobo. Bobo turned out to be a sad-faced clown who refused to speak. It was clear he planned to teach entirely in pantomime. As he silently motioned for me to find my seat, I resigned myself to a bad experience. So imagine my surprise when this crazy class turned out to be invaluable to me! Although I didn’t hear a single spoken word for five hours, Bobo somehow managed to express his ideas clearly and in an engaging way. By the end of the first session, he had expertly taught me how to make balloon animals, apply makeup, juggle beanbags, and develop my own personal clown character. As I eagerly await the second session, "All About Unicycles," I will begin clowning around as a volunteer at a local children’s hospital. Which sentence best summarizes the author’s point of view about the clown class? Choices: [(a) Bobo uses eccentric teaching methods to efficiently teach this course, making him an excellent clowning instructor., (b) Bobo’s sad demeanor and poor pantomime skills make it difficult for his students to learn effective clowning techniques.] Answer index: 0
@@
Subject: Language Arts Skill: compare-two-texts-with-different-genres Description: Read the following text from a diary. This excerpt is about life for Jews in Europe at the beginning of World War II. After May 1940 good times rapidly fled: first the war, then the capitulation, followed by the arrival of the Germans, which is when the sufferings of us Jews really began. Anti-Jewish decrees followed each other in quick succession. Jews must wear a yellow star, Jews must hand in their bicycles, Jews are banned from trams and are forbidden to drive. Jews are only allowed to do their shopping between three and five o’clock and then only in shops which bear the placard "Jewish shop." Jews must be indoors by eight o’clock and cannot even sit in their own gardens after that hour. Jews are forbidden to visit theaters, cinemas, and other places of entertainment. From Anne Frank, "Anne Frank: The Diary of a Young Girl". Copyright 1952 by Otto H. Frank Based on this excerpt, what can you conclude about life for Jews in Europe at the beginning of World War II? Choices: [(a) Jewish people could no longer afford to buy theater tickets., (b) Jewish people could not associate with one another., (c) Jewish people had many of their rights taken away.] Answer index: 2
@@
Subject: Language Arts Skill: read-and-understand-informational-passages Description: Review the **seventh paragraph**. title: Meditating in School This is where the need for good, sound research comes in. Past studies showing that mindfulness helps medical patients are not enough. To really determine whether mindfulness helps improve grades, research needs to focus specifically on mindfulness in schools—and what impact, if any, it has. That way, educators can make informed decisions about where to put their resources. Until then, whether mindfulness is included in classroom programming will remain the choice of individual teachers. Based on the text, what still needs to be determined about mindfulness in schools? Choices: [(a) whether meditation affects different students in different ways, (b) how to raise the funds needed to teach meditation in schools, (c) whether mindfulness meditation has a positive impact in a school setting] Answer index: 2
@@
Subject: Language Arts Skill: choose-evidence-to-support-a-claim Description: Read the claim below. Gardeners should never use pesticides on their plants. Select the piece of evidence that best supports this claim. Choices: [(a) Organic pesticides tend to cost more than conventional ones, and they may be less effective., (b) Pets and children often play in and around gardens, where they can be exposed to pesticides.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-the-giver-part-1 Description: Reread the following passage from The Giver. In this excerpt, Jonas, the main character, recalls a time when he saw an unidentified aircraft. Frightened was the way [Jonas] had felt a year ago when an unidentified aircraft had overflown the community twice. He had seen it both times. Squinting toward the sky, he had seen the sleek jet, almost a blur at its high speed, go past, and a second later heard the blast of sound that followed. Then one more time, a moment later, from the opposite direction, the same plane. From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry Look at the text in **bold**. How does this paragraph contribute to the meaning or tone of the passage? Choices: [(a) It creates an unsettling tone., (b) It creates a ridiculous tone., (c) It creates a playful tone., (d) It creates a serene tone.] Answer index: 0
Subject: Language Arts Skill: use-personification Description: Complete the sentence so that it uses personification. The sorceress peered into the mirror, noticing how the amulet ___ her neck. Choices: [(a) barely clasped around, (b) gently embraced] Answer index: 1
Subject: Language Arts Skill: identify-plagiarism Description: Compare the student text with the source. Source: Carolyn Cinami DeCristofano, "A Black Hole Is Not a Hole", page 5. Published by Charlesbridge, 2012: Nothing can out-tug a black hole. No army of tow trucks, no convoy of supersized earth haulers, no fleet of giant rocket engines. Student text: The pull of a black hole is so strong that nothing can pull stronger, not even a "fleet of giant rocket engines" (DeCristofano 5). Is the student text plagiarized? Choices: [(a) No, it is not plagiarized., (b) Yes, because it fails to use quotation marks.] Answer index: 0
Subject: Language Arts Skill: identify-counterclaims Description: Consider this claim: Poems that rhyme are better than poems that do not rhyme. Now consider how someone might argue against this. Which sentence presents a counterclaim to the above claim? Choices: [(a) "Beowulf" is one of the oldest and longest poems in existence., (b) Poetry that doesn’t rhyme allows for more freedom of expression.] Answer index: 1
@@
Subject: Language Arts Skill: trace-an-argument Description: Read the text. title: Early to Bed and Early to Rise Whether you’re naturally a night owl or an early bird is mostly based on your body’s chronotype, which Merriam-Webster’s online dictionary defines as "the internal circadian rhythm or body clock of an individual that influences the cycle of sleep and activity in a 24-hour period." In short, we are born wired one way or the other. However, it is possible to change your habits. And given the benefits of being an early riser, it may be worth the effort. What is the author’s main claim or argument? Choices: [(a) There are just as many benefits to being a night owl as an early riser., (b) It’s important to change how society perceives night owls., (c) It is better to be an early riser than a night owl.] Answer index: 2
Subject: Language Arts Skill: identify-active-and-passive-voice Description: Which sentence is in passive voice? Choices: [(a) On April 18, 1906, an earthquake and its fiery aftermath devastated the city of San Francisco., (b) On April 18, 1906, San Francisco was devastated by an earthquake and its fiery aftermath.] Answer index: 1
@@
Subject: Language Arts Skill: is-the-pronoun-reflexive-or-intensive Description: Is the pronoun in bold reflexive or intensive? Bryce uses an elaborate system of notes to remind **himself** of what he has left to do. Choices: [(a) intensive, (b) reflexive] Answer index: 1
Subject: Language Arts Skill: interpret-figures-of-speech Description: What does the **metaphor** in this text suggest? When Bobby lost his job, he was devastated. The only light in the sea of darkness was the prospect of pursuing a new career. Choices: [(a) There was a benefit to Bobby’s job loss., (b) Having to pursue a new career was the worst part of Bobby’s job loss.] Answer index: 0
Subject: Language Arts Skill: identify-the-verb-mood Description: Which verb mood is used in the **bold** part of the sentence? To help maintain your energy and focus throughout the day, **eat a healthy breakfast**. Choices: [(a) imperative, (b) subjunctive] Answer index: 0
Subject: Language Arts Skill: classify-figures-of-speech Description: Which figure of speech is used in this text? The water made a sound **like kittens lapping**. —Marjorie Kinnan Rawlings, "The Yearling" Choices: [(a) metaphor, (b) simile] Answer index: 1
@@
Subject: Language Arts Skill: vocabulary-review-trace-an-argument Description: Look at the word **entrenched** as used in the passage "The Most Important Meal of the Day?" An Australian study published in 2019 collected health data for over twenty-eight years from multiple countries. The results showed few differences between breakfast skippers and breakfast eaters. The study concluded that "breakfast is not the most important time of the day to eat, even though that belief is really **entrenched** in our society and around the world." What does **entrenched** mean? Choices: [(a) highly debatable, (b) quite rare, (c) firmly established] Answer index: 2
@@
Subject: Language Arts Skill: analyze-passages-from-the-outsiders-part-2 Description: Review the passage. Tim Shepard and company were already waiting when we arrived at the vacant lot, along with a gang from Brumly, one of the suburbs. Tim was a lean, catlike eighteen-year-old who looked like the model JD you see in movies and magazines. He had the right curly black hair, smoldering dark eyes, and a long scar from temple to chin where a tramp had belted him with a broken pop bottle. He had a tough, hard look to him, and his nose had been broken twice. Like Dally’s, his smile was grim and bitter. He was one of those who enjoy being a hood. The rest of his bunch were the same way. The boys from Brumly, too. Young hoods—who would grow up to be old hoods. I’d never thought about it before, but they’d just get worse as they got older, not better. I looked at Darry. He wasn’t going to be any hood when he got old. He was going to get somewhere. Living the way we do would only make him more determined to get somewhere. That’s why he’s better than the rest of us, I thought. He’s going somewhere. And I was going to be like him. I wasn’t going to live in a lousy neighborhood all my life. From S. E. Hinton, The Outsiders. Copyright 1967 by S. E. Hinton Look at the text in **bold**. What does this figurative language suggest? Choices: [(a) Tim fights for survival every day., (b) Tim has always been cruel to Ponyboy., (c) Tim often does not have enough to eat., (d) Tim attempts to stay hidden at all costs.] Answer index: 0
@@
Subject: Language Arts Skill: identify-the-authors-purpose Description: Review the text. Good study habits train your brain for success. Applying effective learning methods will not only improve your grades but, more importantly, increase your long-term knowledge. Cramming at the last minute is ineffective. Instead, try to create a study routine. Schedule specific blocks of time to study every day. Next, find the best environment that helps you concentrate. Some students do better with the silence of a library, while others need the background noise or music of a public space. Now that you have a time and a place, you need a strategy. Look at your tasks and make a list. Setting small, manageable goals for your study session will help you work efficiently. Do the most difficult assignments first, while your mind is fresh. Finally, do not be afraid to take limited breaks to restore your motivation. Which best shows that the author’s purpose is "to explain some productive study techniques to readers"? Choices: [(a) The author argues that students need to work in a place with "background noise" in order to study efficiently., (b) The author outlines a series of best practices to follow to improve study habits.] Answer index: 1
Subject: Language Arts Skill: is-it-a-direct-object-or-an-indirect-object Description: Is the phrase in bold a direct object or an indirect object? A guard denied **the tourists** entrance to the palace. Choices: [(a) direct object, (b) indirect object] Answer index: 1
Subject: Language Arts Skill: decide-whether-ellipses-are-used-appropriately Description: Read this text. On May 25, 1977, "Star Wars" opened in theaters and wowed audiences with groundbreaking special effects. Is this an appropriate use of an ellipsis? Choices: [(a) yes, (b) no, because the text with an ellipsis is no longer grammatically correct] Answer index: 1
Subject: Language Arts Skill: compound-subjects-and-objects-with-pronouns Description: Select the correct pronoun to complete the sentence. Mr. Joyce rarely worries about ___ and Farid, because they’re both so responsible. Choices: [(a) she, (b) her] Answer index: 1
Subject: Language Arts Skill: identify-vague-pronoun-references Description: Which of the following contains a vague pronoun reference? Choices: [(a) Bella and her sister posed for a photograph, but when the flash went off, she blinked., (b) Bella and her sister posed for a photograph, but when the flash went off, Bella blinked.] Answer index: 0
@@
Subject: Language Arts Skill: what-does-the-punctuation-suggest Description: What does the following sentence suggest? Ferries depart from Battery Park in New York and Liberty State Park in New Jersey for the museum on Ellis Island, which commemorates the millions of immigrants who entered the United States through Ellis Island. Choices: [(a) There is only one museum on Ellis Island., (b) There is more than one museum on Ellis Island.] Answer index: 0
@@
Subject: Language Arts Skill: suggest-appropriate-revisions Description: Read the following text from a student essay. How can the writer best improve his or her **word choice**? I saw a movie called "Blob Invasion". It was really good. It was about a weird guy who was into science. He worked on bugs and small animals. He made a machine that did things to the bugs and the animals. They became big weird things. One day he transformed himself by accident. He became a big blob! He went around doing bad things to the city. In the end, a boy figured out how to change the blob back into a man. I would recommend this movie to everyone who likes fun science movies. Choices: [(a) by reducing repetitive language, (b) by using more specific language] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-the-giver-part-2 Description: Review the passage. "Father? Mother?" Jonas asked tentatively after the evening meal. "I have a question I want to ask you." From Lois Lowry, The Giver. Copyright 1993 by Lois Lowry Based on the passage, how does Jonas most likely feel about his family? Choices: [(a) He finds comfort in the regimented order provided by his family., (b) He wonders why his family’s beliefs are different from the community’s., (c) He feels grateful to have a family that openly discusses tough issues., (d) He is beginning to question his parents and the community.] Answer index: 3
@@
Subject: Language Arts Skill: identify-thesis-statements Description: Which is a thesis statement? Choices: [(a) School-wide Wi-Fi access is critical to building academic success because it increases communication, allows for more interactive lessons, and expands research opportunities., (b) What effect does having Wi-Fi access in public school classrooms and libraries have on how students perform academically and how well they pay attention in class?] Answer index: 0
@@
Subject: Language Arts Skill: classify-logical-fallacies Description: Which logical fallacy is used in the text? In this election, we’ve seen the highest voter turnout ever recorded. If you have any doubts about Ryan Mueller’s qualifications, just look at how many people have come out to vote for him. Choices: [(a) bandwagon fallacy: the assumption that the popular choice is automatically correct, (b) ad hominem: a personal attack against one’s opponent] Answer index: 0